Mohave Corp. makes several varieties of beach umbrellas and accessories. It has been approached by a company called Lost Mine Industries about producing a special order for a custom umbrella called the Ultimate Shade (US). The special-order umbrellas with the Lost Mine Company logo would be distributed to participants at an upcoming convention sponsored by Lost Mine.

Lost Mine has offered to buy 2,400 of the US umbrellas at a price of $20 each. Mohave currently has the excess capacity necessary to accept the offer. The following information is related to the production of the US umbrella:   

Direct materials $ 9.00
Direct labor 5.00
Variable manufacturing overhead 5.50
Fixed manufacturing overhead 2.50
Total cost $ 22.00
Regular sales price $ 28.00

Required:
1.
Compute the incremental profit (or loss) from accepting the special order.

Profit (or Loss): by .

2. Should Mohave accept the special order?

No
Yes



3. Suppose that the special order had been to purchase 2,900 umbrellas for $17.00 each. Recompute the incremental profit (or loss) from accepting the special order under this scenario.

Profit (or Loss): by .

4. Assume that Mohave is operating at full capacity. Calculate the special-order price per unit at which Mohave would be indifferent between accepting or rejecting the special order.

Special Order Price: /unit
Answer #1

1.

Special order price = $20 per unit

Variable cost per unit = Direct material + Direct labor + Variable manufacturing overhead

= 9 + 5 + 5.50

= $19.50

Profit per unit in special order = Special order price - Variable cost per unit

= 20 - 19.50

= $0.50

Special order size = 2,400 units

Hence, incremental profit from accepting the special order = 2,400 x 0.50

= $1,200

2.

Special order should be accepted since it will provide an additional income of $1,200

3.

Special order price = $17 per unit

Variable cost per unit = Direct material + Direct labor + Variable manufacturing overhead

= 9 + 5 + 5.50

= $19.50

Loss per unit in special order = Special order price - Variable cost per unit

= 17 - 19.50

= - $2.50

Special order size = 2,900 units

Hence, incremental loss from accepting the special order = 2,900 x - 2.50

= - $7,250

4.

If Mohave is operating at full capacity,  the special-order price per unit at which Mohave would be indifferent between accepting or rejecting the special order must cover the total cost of producing a unit.

Hence, at the full capacity, special order should not be accepted at a price lower than total cost i.e. $22 per unit.

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.